# Towards a generic absoluteness theorem for Chang models

### v2

@@ 10.1. Resurrection of derived models
    Suppose for some $\lambda$, $F\in V^{*}$ is a $(j(\kappa),\lambda)$-extender and $i:V^{*}\rightarrow M=\operatorname{Ult}(V^{*},F)$ is the ultrapower embedding given by $F$. Let $G\subseteq\operatorname{Col}(\omega,{<}j(\kappa))$ be $V^{*}$-generic and let, in $V^{*}[G]$, $\mathcal{F}\subseteq{\wp}_{\omega_{1}}(\Gamma^{\infty}_{G})$ be the filter given by $A\in\mathcal{F}$ if and only if for any $M[G]$-generic $H\subseteq\operatorname{Col}(\omega,{<}[j(\kappa),i(j(\kappa))))$ [^56] The conditions in this poset are finite functions $q:[j(\kappa),i(j(\kappa)))\times\omega\rightarrow i(j(\kappa))$ such that $q(\alpha,n)<\alpha$ for every $(\alpha,n)\in dom(q)$., letting $i^{+}:V^{*}[G]\rightarrow M[G*H]$ be the lift-up of $i$, $i^{+}\,"\Gamma^{\infty}_{G}\in i^{+}(A)$. Then letting $\mathcal{F}^{\prime}=\mathcal{F}\cap L(\Gamma^{\infty}_{G},\mathbb{R}_{G})[\mathcal{F}]$,
@@
      $L(\Gamma^{\infty}_{G},\mathbb{R}_{G})[\mathcal{F}]\models``\mathcal{F}^{\prime}$ is a normal, fine ultrafilter on ${\wp}_{\omega_{1}}(\Gamma^{\infty}_{G})"$ and
@@
By elementarity we have that the following holds in $V$.
@@
    Suppose for some $\lambda$, $F\in V$ is a $(\kappa,\lambda)$-extender and $i:V\rightarrow M=\operatorname{Ult}(V,F)$ is the ultrapower embedding given by $F$. Let $G\subseteq\operatorname{Col}(\omega,{<}\kappa)$ be $V$-generic and let, in $V[G]$, $\mathcal{F}\subseteq{\wp}_{\omega_{1}}(\Gamma^{\infty}_{G})$ be the filter given by $A\in\mathcal{F}$ if and only if for any $M[G]$-generic $H\subseteq\operatorname{Col}(\omega,{<}[\kappa,i(\kappa)))$, letting $i^{+}:V[G]\rightarrow M[G*H]$ be the lift-up of $i$, $i^{+}\,"\Gamma^{\infty}_{G}\in i^{+}(A)$. Then letting $\mathcal{F}^{\prime}=\mathcal{F}\cap L(\Gamma^{\infty}_{G},\mathbb{R}_{G})[\mathcal{F}]$,
@@
      $L(\Gamma^{\infty}_{G},\mathbb{R}_{G})[\mathcal{F}]\models``\mathcal{F}^{\prime}$ is a normal, fine ultrafilter on ${\wp}_{\omega_{1}}(\Gamma^{\infty}_{G})"$ and
@@
Also, notice that letting $\delta={\sf{ehst}}(\kappa)$ and letting $k:Ult(V,F_{\delta})\rightarrow Ult(V,F)$ be the canonical factor embedding given by $k([a,f]_{F_{\delta}})=[a,f]_{F}$, ${\rm cp}(k)=\delta$ [^57] This does not mean that we have set $\delta={\sf{ehst}}(\kappa)$..
Suppose now that $h$ is $V[g]$-generic for a poset $\mathbb{P}\in V_{j(\kappa)}[g]$. Let $\lambda<\chi_{0}<\chi_{1}<j(\kappa)$ be such that for $i\in 2$, $\chi_{i}$ is an inaccessible limit of Woodin cardinals of $V[g]$, and $\lambda>\delta$ is an inaccessible cardinal such that $\mathbb{P}\in V_{\lambda}[g]$. Let $W\in V_{\chi_{0}}$ be a $(\chi_{1},\lambda^{+},\lambda)$-saturated set of extenders that is infinitely tamed above $\lambda$, and let $X\prec V_{\chi+1}$ be such that $V_{\delta+1}\cup\{\lambda,F_{\chi_{0}},W\}\in X$ and ${|X|}={|V_{\delta+1}|}$.
We now use Theorem 6.6 in $V^{*}[g*h]$ [^58] Also, see Remark 6.7.. Let $\iota=\omega_{1}^{V[g*h]}$ and let $k\subseteq\operatorname{Col}(\alpha,{|\Gamma^{\infty}_{g*h}|})$ be $V[g*h]$-generic. We have that since $k$ is generic for a countably closed poset, $\Gamma^{\infty}_{g*h*k}=\Gamma^{\infty}_{g*h}$. Let
@@
be the underlying sequence of ${\sf{DM}}$, $M_{\iota}$ be the direct limit of ${\sf{dm}}$ and $\sigma_{\iota}:M_{\iota}\rightarrow U_{\chi}$ be the canonical realizable embedding.
@@
It follows from Theorem 6.6 that there is an $M_{\iota}$-generic $m\subseteq\operatorname{Col}(\omega,{<}\iota)$ such that $m\in V[g*h*k]$ and $\Gamma^{\infty}_{g*h}=(Hom^{*})^{M_{\iota}[m]}$.
Let
@@
    for $i\leq\iota$, $hom_{i}=(Hom^{*})^{M_{i}[m_{i}]}$ [^59] We thus have that $hom_{i}=(\Gamma^{\infty})^{M_{i}[m_{i}]}$.,
@@
    for $i\leq\iota$, let $\mathcal{F}_{i}\subseteq({\wp}_{\omega_{1}}(Hom^{*}))^{M_{i}[m_{i}]}$ be the filter derived from $\sigma_{i}$. More precisely, $A\in\mathcal{F}_{i}$ if and only if for every $\operatorname{Ult}(M_{i},U_{i})[m_{i}]$-generic
@@
    $\tau_{i}^{+}\,"hom_{i}\in\tau_{i}^{+}(A)$ where $\tau^{+}_{i}:M_{i}[m_{i}]\rightarrow\operatorname{Ult}(M_{i},U_{i})[m_{i}*n]$ is the lift-up of $\tau_{i}$,
@@
Because for $i\leq\iota$, $U_{i}$ is a sub-extender of $E_{i}$, there is $\tau^{\prime}_{i}:\operatorname{Ult}(M_{i},U_{i})\rightarrow M_{i+1}$ with the property that $\tau_{i}^{\prime}\circ\tau_{i}=\sigma_{i,i+1}$. Moreover, ${\rm cp}(\tau_{i}^{\prime})=\lambda_{i}$ [^60] This fact is key in (5.3).. Let now for $i<i^{\prime}\leq\iota$, $\sigma_{i,i^{\prime}}^{+}:M_{i}[m_{i}]\rightarrow M_{i^{\prime}}[m_{i^{\prime}}]$ be the lift-up of $\sigma_{i,i^{\prime}}$. We now have that
@@
    if $A\in\mathcal{F}_{i}$ then $\sigma^{+}_{i,i+1}\,"hom_{i}\in\sigma^{+}_{i,i+1}(A)$ [^61] This is because we can lift $\tau_{i}^{\prime}$ to $\tau_{i}^{\prime+}:\operatorname{Ult}(M_{i},U_{i})[m\cap\operatorname{Col}(\omega,{<}\lambda_{i})]\rightarrow M_{i+1}[m_{i+1}]$.,
@@
    $A\in\mathcal{F}_{\iota}$ then for some $i<\iota$, $\{\sigma^{+}_{i,\iota}\,"hom_{i}\mid i<\iota\}\subseteq A$ [^62] This follows from (5.1), (5.3) and the fact that $\iota$ is a limit ordinal..
It then follows from (4.1) and (5.4)
@@
    in $V[g*h*H]$, $\mathcal{F}_{\iota}^{\prime}=\mathcal{C}^{\infty}\cap M_{\alpha}[m]$.
@@
Notice next that $L(\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h}]=L(\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h*k}]$. This is because if $A\in V[g*h]$, $A\subseteq{\wp}_{\iota}(\Gamma^{\infty}_{g*h})$ and $A$ is stationary in $V[g*h]$ then $A$ is stationary in $V[g*h*k]$ [^63] This follows from the fact that countably closed posets are proper.. Therefore, we have that
@@
    in $V[g*h]$, letting $\mathcal{C}^{\prime}=\mathcal{C}^{\infty}\cap L(\Gamma^{\infty},\mathbb{R})[\mathcal{C}^{\infty}]$, we have that
@@
such that $j(\mathcal{C}^{\infty}_{g*h})=\mathcal{C}^{\infty}_{g*h*h^{\prime}}$. What we have shown is the following (we use the objects introduced above). Using the above notation, given $d\in L(\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h}]$ and $i<\iota$ such that $d\in\operatorname{rng}(\sigma_{i,\iota}^{+})$, we let $d_{i}$ be the $\sigma_{i,\iota}$-preimage of $d$. Given sets $a,b,c$, we let $((a,b)[c])^{\#}$ be the sharp of $L(a,b)[c]$ [^64] The language of such a sharp contains constants for the members of $a$ and $b$ and another constant symbol for $c\cap L(a,b)[c]$.. Let $\mathcal{C}^{\prime}_{g*h}=\mathcal{C}^{\infty}_{g*h}\cap L(\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h}]$.
@@
    $(\phi,A,\mathcal{C}^{\prime}_{g*h},c_{0},...,c_{n})\in(\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h}])^{\#}$
@@
    $(\phi,A,\mathcal{C}^{\prime}_{g*h},c_{0},...,c_{n})\in(\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h}])^{\#}$
@@
    $(\phi,A_{h^{\prime}},\mathcal{C}^{\prime}_{g*h*h^{\prime}},c_{0},...,c_{n})\in(\Gamma^{\infty}_{g*h*h^{\prime}},\mathbb{R}_{g*h*h^{\prime}})[\mathcal{C}^{\infty}_{g*h*h^{\prime}}])^{\#}$

### v3

@@ 10.1. Resurrection of derived models
    Suppose for some $\lambda$, $F\in V^{*}$ is a $(j(\kappa),\lambda)$-extender and $i:V^{*}\rightarrow M=Ult(V^{*},F)$ is the ultrapower embedding given by $F$. Let $G\subseteq\operatorname{Col}(\omega,{<}j(\kappa))$ be $V^{*}$-generic and let, in $V^{*}[G]$, $\mathcal{F}\subseteq{\wp}_{\omega_{1}}(\Gamma^{\infty}_{G})$ be the filter given by $A\in\mathcal{F}$ if and only if for any $M[G]$-generic $H\subseteq\operatorname{Col}(\omega,{<}[j(\kappa),i(j(\kappa))))$ [^60] The conditions in this poset are finite functions $q:[j(\kappa),i(j(\kappa)))\times\omega\rightarrow i(j(\kappa))$ such that $q(\alpha,n)<\alpha$ for every $(\alpha,n)\in dom(q)$., letting $i^{+}:V^{*}[G]\rightarrow M[G*H]$ be the lift-up of $i$, $i^{+}\,"\Gamma^{\infty}_{G}\in i^{+}(A)$. Then, letting
    $\mathcal{F}^{\prime}=\mathcal{F}\cap L(\Gamma^{\infty}_{G},\mathbb{R}_{G})[\mathcal{F}],$
@@
      $L(\Gamma^{\infty}_{G},\mathbb{R}_{G})[\mathcal{F}]\models``\mathcal{F}^{\prime}$ is a normal, fine ultrafilter on ${\wp}_{\omega_{1}}(\Gamma^{\infty}_{G})"$, and
@@
By elementarity, we have that the following holds in $V$.
@@
    Suppose for some $\lambda$, $F\in V$ is a $(\kappa,\lambda)$-extender and $i:V\rightarrow M=Ult(V,F)$ is the ultrapower embedding given by $F$. Let $G\subseteq\operatorname{Col}(\omega,{<}\kappa)$ be $V$-generic and let, in $V[G]$, $\mathcal{F}\subseteq{\wp}_{\omega_{1}}(\Gamma^{\infty}_{G})$ be the filter given by $A\in\mathcal{F}$ if and only if for any $M[G]$-generic $H\subseteq\operatorname{Col}(\omega,{<}[\kappa,i(\kappa)))$, letting $i^{+}:V[G]\rightarrow M[G*H]$ be the lift-up of $i$, $i^{+}\,"\Gamma^{\infty}_{G}\in i^{+}(A)$. Then, letting
    $\mathcal{F}^{\prime}=\mathcal{F}\cap L(\Gamma^{\infty}_{G},\mathbb{R}_{G})[\mathcal{F}],$
@@
      $L(\Gamma^{\infty}_{G},\mathbb{R}_{G})[\mathcal{F}]\models``\mathcal{F}^{\prime}$ is a normal, fine ultrafilter on ${\wp}_{\omega_{1}}(\Gamma^{\infty}_{G})"$, and
@@
Also, notice that letting $\delta={\sf{ehst}}(\kappa)$ and letting $k:Ult(V,F_{\delta})\rightarrow Ult(V,F)$ be the canonical factor embedding given by $k([a,f]_{F_{\delta}})=[a,f]_{F}$, ${\rm cp}(k)=\delta$.
Suppose now that $h$ is $V[g]$-generic for a poset $\mathbb{P}\in V_{j(\kappa)}[g]$. Let $\lambda<\chi_{0}<\chi_{1}<j(\kappa)$ be such that for $i\in 2$, $\chi_{i}$ is an inaccessible limit of Woodin cardinals of $V[g]$, and $\lambda>\delta$ is an inaccessible cardinal such that $\mathbb{P}\in V_{\lambda}[g]$. Let $W\in V_{\chi_{0}}$ be a $(\chi_{1},\lambda^{+},\lambda)$-saturated set of measures that is infinitely tamed above $\lambda$, and let $X\prec V_{\chi_{1}+1}$ be such that $V_{\delta+1}\cup\{\lambda,F_{\chi_{0}},W\}\in X$ and ${|X|}={|V_{\delta+1}|}$.
We now use Theorem 6.6 in $V^{*}[g*h]$ [^61] Also, see Remark 6.7.. Let $\iota=\omega_{1}^{V[g*h]}$ and let $k\subseteq\operatorname{Col}(\iota,{|\Gamma^{\infty}_{g*h}|})$ be $V[g*h]$-generic. We have that since $k$ is generic for a countably closed poset, $\Gamma^{\infty}_{g*h*k}=\Gamma^{\infty}_{g*h}$. Let
@@
be the underlying sequence of $D$, $M_{\iota}$ be the direct limit of ${\sf{dm}}$ and $\sigma_{\iota}:M_{\iota}\rightarrow U_{\chi}$ be the canonical realizability embedding.
@@
It follows from Theorem 6.6 that there is an $M_{\iota}$-generic $m\subseteq\operatorname{Col}(\omega,{<}\iota)$ such that $m\in V[g*h*k]$ and $\Gamma^{\infty}_{g*h}=(Hom^{*})^{M_{\iota}[m]}$. Let
@@
    for $i\leq\iota$, $hom_{i}=(Hom^{*})^{M_{i}[m_{i}]}$, [^62] We thus have that $hom_{i}=(\Gamma^{\infty})^{M_{i}[m_{i}]}$.
@@
    for $i\leq\iota$, let $\mathcal{F}_{i}\subseteq({\wp}_{\omega_{1}}(Hom^{*}))^{M_{i}[m_{i}]}$ be the filter derived from $\sigma_{i}$. More precisely, $A\in\mathcal{F}_{i}$ if and only if for every $Ult(M_{i},U_{i})[m_{i}]$-generic
@@
    $\tau_{i}^{+}\,"hom_{i}\in\tau_{i}^{+}(A)$ where $\tau^{+}_{i}:M_{i}[m_{i}]\rightarrow Ult(M_{i},U_{i})[m_{i}*n]$ is the lift-up of $\tau_{i}$,
@@
Because $U_{i}$ is a sub-extender of $E_{i}$ for $i\leq\iota$, there is $\tau^{\prime}_{i}:Ult(M_{i},U_{i})\rightarrow M_{i+1}$ with the property that $\tau_{i}^{\prime}\circ\tau_{i}=\sigma_{i,i+1}$. Moreover, ${\rm cp}(\tau_{i}^{\prime})=\lambda_{i}$. [^63] This fact is key in (5.3). Let now for $i<i^{\prime}\leq\iota$, $\sigma_{i,i^{\prime}}^{+}:M_{i}[m_{i}]\rightarrow M_{i^{\prime}}[m_{i^{\prime}}]$ be the lift-up of $\sigma_{i,i^{\prime}}$. We now have that
@@
    if $A\in\mathcal{F}_{i}$ then $\sigma^{+}_{i,i+1}\,"hom_{i}\in\sigma^{+}_{i,i+1}(A)$, [^64] This is because we can lift $\tau_{i}^{\prime}$ to $\tau_{i}^{\prime+}:Ult(M_{i},U_{i})[m\cap\operatorname{Col}(\omega,{<}\lambda_{i})]\rightarrow M_{i+1}[m_{i+1}]$.
@@
    if $A\in\mathcal{F}_{\iota}$ then for some $i<\iota$, $\{\sigma^{+}_{i,\iota}\,"hom_{i}\mid i<\iota\}\subseteq A$ [^65] This follows from (5.1), (5.3) and the fact that $\iota$ is a limit ordinal.
It then follows from (4.1) and (5.4) that
@@
    in $V[g*h*H]$, $\mathcal{F}_{\iota}^{\prime}=\mathcal{C}^{\infty}\cap M_{\iota}[m]$.
@@
Notice next that $L(\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h}]=L(\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h*k}]$. This is because if $A\in V[g*h]$, $A\subseteq{\wp}_{\iota}(\Gamma^{\infty}_{g*h})$ and $A$ is stationary in $V[g*h]$ then $A$ is stationary in $V[g*h*k]$. [^66] This follows from the fact that countably closed posets are proper. Therefore, we have that
@@
    in $V[g*h]$, letting $\mathcal{C}^{\prime}=\mathcal{C}^{\infty}\cap L(\Gamma^{\infty},\mathbb{R})[\mathcal{C}^{\infty}]$,
@@
such that $j(\mathcal{C}^{\infty}_{g*h})=\mathcal{C}^{\infty}_{g*h*h^{\prime}}$. What we have shown is the following (we use the objects introduced above). Using the above notation, given $d\in L(\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h}]$ and $i<\iota$ such that $d\in\operatorname{rng}(\sigma_{i,\iota}^{+})$, we let $d_{i}$ be the $\sigma_{i,\iota}$-preimage of $d$. Given sets $a,b,c$, we let $((a,b)[c])^{\#}$ be the sharp of $L(a,b)[c]$. [^67] The language of such a sharp contains constants for the members of $a$ and $b$ and another constant symbol for $c\cap L(a,b)[c]$. Let $\mathcal{C}^{\prime}_{g*h}=\mathcal{C}^{\infty}_{g*h}\cap L(\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h}]$.
@@
    $(\phi,A,\mathcal{C}^{\prime}_{g*h},c_{0},...,c_{n})\in((\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h}])^{\#}$
@@
    $(\phi,A,\mathcal{C}^{\prime}_{g*h},c_{0},...,c_{n})\in((\Gamma^{\infty}_{g*h},\mathbb{R}_{g*h})[\mathcal{C}^{\infty}_{g*h}])^{\#}$
@@
    $(\phi,A_{h^{\prime}},\mathcal{C}^{\prime}_{g*h*h^{\prime}},c_{0},...,c_{n})\in((\Gamma^{\infty}_{g*h*h^{\prime}},\mathbb{R}_{g*h*h^{\prime}})[\mathcal{C}^{\infty}_{g*h*h^{\prime}}])^{\#}$